# HUNT: High-Speed UAV Navigation and Tracking in Unstructured Environments via Instantaneous Relative Frames

### v1

@@ Abstract
Search and rescue operations require unmanned aerial vehicles to both traverse unknown unstructured environments at high speed and track targets once detected. Achieving both capabilities under degraded sensing and without global localization remains an open challenge. Recent works on relative navigation have shown robust tracking by anchoring planning and control to a visible detected object, but cannot address navigation when no target is in the field of view. We present HUNT (High-speed UAV Navigation and Tracking), a real-time framework that unifies traversal, acquisition, and tracking within a single relative formulation. HUNT defines navigation objectives directly from onboard instantaneous observables such as attitude, altitude, and velocity, enabling reactive high-speed flight during search. Once a target is detected, the same perception–control pipeline transitions seamlessly to tracking. Outdoor experiments in dense forests, container compounds, and search-and-rescue operations with vehicles and mannequins demonstrate robust autonomy where global methods fail. Video: http://bit.ly/4n5Tp5D
@@ III-A2 Reference Generation
At each stage $i\in[0,N)$ of the NMPC prediction horizon of length $N$ and discretization step $\Delta t$, the reference trajectory $\{\bar{\mathbf{x}}_{t+i}\}$ is rolled out by advancing at constant speed $v^{\mathrm{loiter}}$ along $\psi_{t}^{\mathrm{loiter}}$ while holding altitude at $p_{z}^{\mathrm{loiter}}$. Formally,
@@ III-B2 Reference Generation
At each stage $i\in[0,N)$ of the NMPC prediction horizon of length $N$ and discretization step $\Delta t$, the reference trajectory $\{\bar{\mathbf{x}}_{t+i}\}$ is rolled out by holding the desired displacement and orientation fixed relative to the target. Formally,
@@ III-D Motion Planning and Control
Both loitering and tracking are executed through a single unified controller: an NMPC augmented with CBF constraints. At each cycle, the NMPC solves a finite-horizon problem of length $N$, minimizing the deviation between the mode-specific predicted states $\hat{\mathbf{x}}_{t+i}$ and references $\bar{\mathbf{x}}_{t+i}$:
@@ IV-A Experimental Setup
All experiments are conducted with a quadrotor weighing $1.2\penalty 10000\ $\mathrm{k}\mathrm{g}$\$, with a thrust-to-weight ratio of $4$ and a motor span of $25\penalty 10000\ $\mathrm{c}\mathrm{m}$\$. Onboard computation is provided by an NVIDIA Jetson Orin NX, while low-level stabilization runs on a PixRacer Pro. The PixRacer’s IMU and barometer, centrally mounted and coincident with the body frame $\mathcal{B}$, provide attitude, angular velocity, and altitude. Perception uses an Intel RealSense D455 operating at $320\times 240$ fused with monocular disparity from DepthAnythingV2 [24] (ViT-S) accelerated on TensorRT at $19\penalty 10000\ $\mathrm{m}\mathrm{s}$\$. Target detection is performed by a custom-trained YOLOv11 network, following the work [2]. The camera is mounted downward at $45^{\circ}$ when flying above canopy or buildings, and forward-facing when operating at ground level to maximize the number of features in view.

### v2

@@ Abstract
Search and rescue operations require unmanned aerial vehicles to both traverse unknown unstructured environments at high speed and track targets once detected. Achieving both capabilities under degraded sensing and without global localization remains an open challenge. Recent works on relative navigation have shown robust tracking by anchoring planning and control to a visible detected object, but cannot address navigation when no target is in the field of view. We present HUNT (High-speed UAV Navigation and Tracking), a real-time framework that unifies traversal, acquisition, and tracking within a single relative formulation. HUNT defines navigation objectives directly from onboard instantaneous observables such as attitude, altitude, and velocity, enabling reactive high-speed flight during search. Once a target is detected, the same perception–control pipeline transitions seamlessly to tracking. Outdoor experiments in dense forests, container compounds, and search-and-rescue operations with vehicles and mannequins demonstrate robust autonomy where global methods fail. Video: https://youtu.be/YsSflqPPHhs
@@ III-A2 Reference Generation
At each stage $i\in[0,N)$ of the NMPC prediction horizon of length NN and discretization step $\Delta t$, the reference trajectory $\{\bar{\mathbf{x}}_{t+i}\}$ is rolled out by advancing at constant speed $v^{\mathrm{loiter}}$ along $\psi_{t}^{\mathrm{loiter}}$ while holding altitude at $p_{z}^{\mathrm{loiter}}$. Formally,
@@ III-B2 Reference Generation
At each stage $i\in[0,N)$ of the NMPC prediction horizon of length NN and discretization step $\Delta t$, the reference trajectory $\{\bar{\mathbf{x}}_{t+i}\}$ is rolled out by holding the desired displacement and orientation fixed relative to the target. Formally,
@@ III-D Motion Planning and Control
Both loitering and tracking are executed through a single unified controller: an NMPC augmented with CBF constraints. At each cycle, the NMPC solves a finite-horizon problem of length NN, minimizing the deviation between the mode-specific predicted states $\hat{\mathbf{x}}_{t+i}$ and references $\bar{\mathbf{x}}_{t+i}$:
@@ IV-A Experimental Setup
All experiments are conducted with a quadrotor weighing $1.2\penalty 10000\ $\mathrm{k}\mathrm{g}$\$, with a thrust-to-weight ratio of 44 and a motor span of $25\penalty 10000\ $\mathrm{c}\mathrm{m}$\$. Onboard computation is provided by an NVIDIA Jetson Orin NX, while low-level stabilization runs on a PixRacer Pro. The PixRacer’s IMU and barometer, centrally mounted and coincident with the body frame $\mathcal{B}$, provide attitude, angular velocity, and altitude. Perception uses an Intel RealSense D455 operating at $320\times 240$ fused with monocular disparity from DepthAnythingV2 [24] (ViT-S) accelerated on TensorRT at $19\penalty 10000\ $\mathrm{m}\mathrm{s}$\$. Target detection is performed by a custom-trained YOLOv11 network, following the work [2]. The camera is mounted downward at $45^{\circ}$ when flying above canopy or buildings, and forward-facing when operating at ground level to maximize the number of features in view.